# Length-Induced Embedding Collapse in Transformer-based Models

### v1

@@ 2.3 Discussion
Other Components in Transformer. After discussing how self-attention contributes to length collapse, we proceed to examine the influence of other modules in the Transformer, such as multi-head, residual, and FFN. Fortunately, previous work (Wang et al., 2022b) has talked about whether these components can effectively alleviate the low-pass filtering drawbacks. The proof demonstrates that while these components help preserve high-frequency signals, they do not alter the fact that the MSA block, as a whole, functions solely with the representational power of a low-pass filter. Furthermore, the ability of these models to preserve high-frequency signals is solely determined by their internal parameters and architecture, independent of the input text length. As a result, these modules do not impact our analysis of length collapse in practical models.
@@
Difference from Over-Smoothing in Deeper Layers. In previous research (Wang et al., 2022b), it has been noted that a self-attention module acts as a low-pass filter, causing input feature maps to gradually lose high-frequency signals as the model layers go deeper. Furthermore, other studies (Oono & Suzuki, ; Cai & Wang, 2020) indicate that the node features of Graph Convolutional Networks (GCNs) can become exponentially trapped in the null space of the graph Laplacian matrix. The root cause of this phenomenon is that both graph Laplacian matrices and self-attention matrices consistently exhibit a dominant eigenvector, commonly referred to as the DC component. While these studies address over-smoothing in deeper layers, we focus on how the low-pass filtering process changes as the input sequence lengthens, specifically examining over-smoothing in longer sequences.
@@ 3 Mitigating Length Collapse via Temperature Scaling
As discussed in Section 2.2, self-attention matrix perform low-pass filtering, which narrows the filter space embedding model can express. Furthermore, a longer sentence length will cause a more narrow filter space, which results in the embedding model’s failure in long texts. To address the problem, an intuitive idea is to increase the diversity of embeddings for long texts, making them more distinguishable within the space. Inspired by Eqn. 5, we propose a scaling technique, called Temperature Scaling (TempScale), which directly manipulates the attention map by multiplying $\tau_{s}$ by a constant temperature $\tau$ less than 1. This slows down the filtering process by increasing $\sigma_{a}$.
Figure 4: Two extreme cases of TempScale: larger $\tau$ causes uniform matrix elements in self-attention matrix $\bm{A}$, filtering out high-frequency information, while smaller $\tau$ preserves high-frequency details by selecting specific token representations. The darker the color, the higher the attention score.
Figure 5: Visualization of the value of $\sigma_{a}$ across different text length. $\tau$ is defined in Eqn. 6 for scaling. See more details in Appendix E.3
Based on Eqn. 4, a smaller $\tau_{s}$ results in a larger $\sigma_{s}$ which will further result in a larger $\sigma_{a}$ based on Eqn. 3. In other words, the low-pass filtering rate of the attention matrix decreases as $\tau_{s}$ decreases. Inspired by this, TempScale introduces a temperature $\tau$ to re-scale the self-attention matrix $\bm{A}$. Formally, let $\bm{A}=\mathrm{softmax}\left(\frac{\bm{X}\bm{W}_{Q}(\bm{X}\bm{W}_{K})^{\top}}{\sqrt{d}}\right)$ denote a self-attention matrix. To decrease the low-pass filtering rate of $\bm{A}$, we apply a temperature coefficient $\tau$ to the logits before performing the softmax operation. Specifically, for each row $\bm{p}_{i}$ in the attention score matrix $\frac{\bm{X}\bm{W}_{Q}(\bm{X}\bm{W}_{K})^{\top}}{\sqrt{d}}$, we compute the scaled logits by dividing by a temperature $\tau\in(0,1]$, and then apply the softmax function to obtain the attention weights:
@@
where a lower $\tau$ results a smaller $\tau_{s}$ and furthur a smaller rate of low pass filtering.
Intuitive Explanation. We present the effects of temperature scaling on two extreme cases to illustrate how TempScale works. As shown in Figure 5, when scaling matrix $\bm{A}$ with a relatively large $\tau$, the elements in the final matrix $\bm{A}$ can be approximated as nearly equal. In this scenario, matrix $\bm{A}$ applied to $\bm{X}$ filters out all high-frequency information, causing all token embeddings to become identical. Similarly, when scaling with a smaller temperature, matrix $\bm{A}$ no longer acts as a weighted sum of all token representations but instead selects the representation of a particular token, allowing for the retention of more high-frequency information. If we view the attention matrix as an adjacency matrix, a higher temperature leads to a denser graph, facilitating more information exchange between nodes and resulting in the loss of high-frequency information (Oono & Suzuki, ; Cai & Wang, 2020). In contrast, a lower temperature produces a sparser graph, allowing self-attention to preserve more high-frequency information at individual nodes and prevent over-smoothing between nodes.
@@ 4 Experiments
In this section, we first conduct experiments to validate the effectiveness of our TempScale on MTEB (Muennighoff et al., 2023) and LongEmbed. Then we analyze how different tasks can benefit from TempScale and how different $\tau$ affect performance on datasets from LongEmbed (Zhu et al., 2024). Finally, we adaptively set $\tau$ based on text length $n$ to validate our theoretical analysis.
4.1 Can TempScale Benefit Embedding Models?
Experiment Settings. For long context retrieval, we use 4 real-world tasks curated from long-form QA and summarization in LongEmbed (Zhu et al., 2024) to access embedding models’ factuality in short-query and long-document settings. For other embedding tasks, we consider six other tasks, including classification, clustering, summarization, semantic textual similarity (STS), retrieval, and reranking, comprising 36 datasets from MTEB (Muennighoff et al., 2023). To comprehensively evaluate TempScale, we select several representative Transformer-based embedding models, including: (1) ANCE (Xiong et al., 2021); (2) GTR (Ni et al., 2022); (3) GIST (Solatorio, 2024); (4) BGE (Xiao et al., 2023); (5) E5 (Zhu et al., 2024). These models are fine-tuned from various pretrain language models, including BERT (Kenton & Toutanova, 2019), RoBERTa (Liu et al., 2019), and T5 (Chung et al., 2024). More descriptions of the datasets and models can be found in Appendix D. When evaluating the embedding models, we set the same $\tau$ on the softmax function for the attention modules across all layers within the range of $\{0.1,0.5,0.6,0.7,0.8,0.9,1.0\}$. Unless otherwise specified, all temperatures used for TempScale are set to $0.8$. The metrics used for different tasks are consistent with MTEB and can be found in Appendix C.
Table 1: Average of the main metric (see Appendix C) per task on MTEB English subsets and LongEmbd. Relative Improv. means percentage increase over the performance without TempScale and improvements are highlighted with while decreasing values are denoted by .
@@ 4.1 Can TempScale Benefit Embedding Models?
ANCE 55.27 33.04 29.58 66.32 36.87 49.09 34.02 43.45
$+\text{ours}(\tau=0.9)$ 55.37 33.28 29.56 66.47 36.86 49.25 33.93 43.53
Relative Improv. (%) 0.17 0.73 -0.05 0.22 -0.01 0.32 -0.25 0.18
GTR 55.10 38.65 29.67 70.11 44.98 54.23 37.33 47.15
$+\text{ours}(\tau=0.8)$ 55.51 39.52 29.83 70.26 45.61 54.16 37.33 47.46
Relative Improv. (%) 0.73 2.26 0.54 0.21 1.41 -0.13 0.01 0.65
GIST 64.75 44.77 31.14 75.61 52.77 58.55 38.21 52.26
$+\text{ours}(\tau=0.9)$ 65.00 44.64 31.17 75.59 53.41 58.60 38.35 52.39
Relative Improv. (%) 0.38 -0.29 0.09 -0.03 1.21 0.08 0.36 0.26
BGE 64.79 45.80 31.03 75.88 55.29 58.87 37.46 52.73
$+\text{ours}(\tau=0.8)$ 64.89 45.61 31.51 75.68 56.00 58.97 38.35 53.00
Relative Improv. (%) 0.16 -0.42 1.53 -0.26 1.29 0.17 2.40 0.51
@@
E5 61.72 38.82 30.58 71.77 47.22 53.12 56.01 51.32
$+\text{ours}(\tau=0.8)$ 62.15 40.22 31.11 72.17 47.06 53.47 56.88 51.87
Relative Improv. (%) 0.70 3.61 1.74 0.55 -0.33 0.65 1.56 1.07
Avg Improv. (%) 0.43 1.18 0.77 0.14 0.71 0.22 0.82 0.53
Results. We select the optimal temperature $\tau$ for each model based on their performance across all tasks and organize the experimental results as shown in Table 1. The results show that these embedding models can benefit from our proposed method TempScale across various general tasks with an average improvement of 0.53% and across long context retrieval datasets with an average improvement of 0.82%. In these tasks, some datasets, such as those for STS, have an average text length of only around 10 tokens, whereas the texts in LongEmbed generally exceed 1000 tokens. Our method proves effective across both, demonstrating that using TempScale not only prevents long text collapse but also enhances the embeddings for short texts, leading to improved performance in downstream tasks. In addition, larger context window sizes lead to greater performance improvements, with E5 showing the highest improvement of 1.07%. This could be attributed to larger windows providing more and longer data for adjustment, which allows for greater performance improvement. Overall, our method, TempScale, is designed specifically for long texts but is also capable of improving model performance in general embedding tasks.
@@ 4.2 Further Analysis
How do the classification and clustering tasks benefit from TempScale? For data grouping tasks like classification and clustering, in the case of $N$-class tasks, we can think of the model as learning $N$ classification boundaries. The farther the text embedding is from the boundary, the closer the output probability approaches 1 or 0. As shown in Figure 6 (left), we plot the entropy of the model output probabilities across different length intervals. The model outputs higher entropy for longer texts, which may be because the embeddings of longer texts are positioned closer to the center of the space as described in Figure 1(b), resulting in a shorter distance to various classification boundaries. Meanwhile, in Figure 6 (right), we can also observe that accuracy and entropy follow the same trend: the model achieves higher accuracy when it has lower entropy. In other words, the model performs better if the text embeddings are farther from the boundary. After applying TempScale, a decreased entropy will result in increased accuracy. This further supports the relationship between entropy and accuracy in classification tasks. Moreover, if the model exhibits a more severe length collapse phenomenon, meaning a greater performance drop on longer texts, the more performance improvement it experiences after applying TempScale. This suggests that one possible reason TempScale is effective on shorter text datasets is that it adjusts the distribution differences between texts of varying lengths. As shown in Figure 1(c), after applying TempScale, the distribution of text embeddings across different length intervals becomes more uniform, which facilitates the model in learning length-agnostic parameters.
Table 2: Ablation study on LongEmbed. “w/o quer” means no query embedding adjustment, and “w/o doc” means no document embedding adjustment.
(a) QMSum
Model ANCE GTR GIST BGE E5
TempScale 23.18 20.23 17.85 23.34 34.11
w/o query 22.70 20.27 17.37 23.13 33.86
w/o doc 21.78 19.42 16.83 22.89 31.58
Raw 21.89 19.54 16.61 22.64 31.27
(b) SummScreenFD
Model ANCE GTR GIST BGE E5
TempScale 55.42 62.46 55.55 58.93 93.48
w/o query 53.11 61.91 55.48 57.15 93.57
w/o doc 55.02 62.47 59.12 60.14 93.75
Raw 53.90 61.81 58.58 57.74 93.87
Figure 7: Results on SummScreenFD.
How do the text matching tasks benefit from TempScale? Text matching tasks, like retrieval and STS, compute relevance scores by embedding two texts separately. We conduct an ablation study on LongEmbed to explore how TempScale affects each part of the text. As shown in Table 2(a), the model’s performance under the “w/o query” setting is generally better than that under the “w/o doc” setting. In contrast, Table 2(b) shows the opposite phenomenon. This difference arises from the query length distribution; QMSum has shorter queries, while SummScreenFD features longer ones. Therefore, SummScreenFD can gain more improvements through TempScaling on longer queries. To verify this explanation, we calculate average $\text{NDCG}@100$ for the longest and shortest $20\%$ of queries on SummScreenFD, as shown in Figure 7. The experimental results demonstrate that a lower temperature brings increasing performance improvements for long queries, further confirming the effectiveness of our method on long texts. Similar phenomena on more models are provided in Appendix E.5.
In addition to improvements on longer queries, we are also interested in the performance changes on longer documents. Since most documents in LongEmbed are very long and are truncated due to the context window size, we select NFCorpus and SciFact to investigate the impact. As shown in Table 3(b), we record the average ranking position of the longest positive documents after applying TempScale at different temperatures, where a lower value indicates the documents are ranked higher. The experimental results show that a lower temperature causes the model to rank relevant long documents higher. This could be because length collapse reduces the distinction between long documents, allowing shorter texts with more contextual representations to be ranked higher. Our method, however, enables long texts to also become more contextual, thereby reducing the bias introduced by length. In summary, TempScale enhances the performance of long documents and long queries by introducing TempScale, which helps the embeddings avoid collapsing into a narrow space.
Figure 8: Results on AmazonPolarity.
(a) NFCorpus
@@
Can the temperature be set based on the length of the text? In the previous experiments, we use the same temperature $\tau$ for scaling all texts in the same task. However, our analysis indicates that texts of different lengths have varying filtering rates, so a natural idea is to use different temperatures for texts of different lengths. As shown in Figure 8, We plot the performance trend for texts under the same settings in Figure 6 as the temperature varies. The results indicate that a higher temperature is optimal for short texts, while a lower temperature is preferable for long texts, as confirmed by the results in Figure 7. More results on other models are provided in Appendix E.6. When performing retrieval tasks, we can also set different temperatures for queries and documents to achieve better performance. As shown in Figure 9, on the QMSum dataset, we can consistently achieve better performance by setting a lower temperature for queries. In the Appendix E.6, we provide more results on SummScreenFD, demonstrating that a lower temperature on documents can also induce a higher performance gain.

### v2

@@ 2.3 Discussion
Other Components in Transformer. After discussing how self-attention contributes to Length Collapse, we further examine the impact of other Transformer components, including multi-head attention, residual connections, and feed-forward networks (FFN). Fortunately, prior work (Wang et al., 2022b) has analyzed whether these modules can alleviate the low-pass filtering effect. The findings show that while they help preserve high-frequency signals, they do not change the fact that the MSA block, as a whole, behaves like a low-pass filter. Moreover, their ability to retain high-frequency information depends entirely on internal parameters and architecture, rather than input length. Therefore, these components do not affect our analysis of Length Collapse in practical Transformer models.
@@
Difference from Over-Smoothing in Deeper Layers. In previous research (Wang et al., 2022b), it has been noted that a self-attention module acts as a low-pass filter, causing input feature maps to gradually lose high-frequency signals as the model layers go deeper. Furthermore, other studies (Oono and Suzuki, 2020; Cai and Wang, 2020; Peng et al., 2024b) indicate that the node features of Graph Convolutional Networks (GCNs) can become exponentially trapped in the null space of the graph Laplacian matrix. The root cause of this phenomenon is that both graph Laplacian matrices and self-attention matrices consistently exhibit a dominant eigenvector, commonly referred to as the DC component. While these studies address over-smoothing in deeper layers, we focus on how the low-pass filtering process changes as the input sequence lengthens, specifically examining over-smoothing in longer sequences.
3 How does Length Collapse Lead to Performance Degradation
Building on our analysis of how the self-attention mechanism contributes to Length Collapse, this section investigates its impact on model performance across the MTEB benchmark. The tasks in MTEB can be broadly categorized into classification/clustering and matching tasks. For classification and clustering, a classifier is trained on text embeddings. In matching tasks, performance is evaluated by computing the similarity between embeddings using cosine similarity or the dot product. These matching tasks can be further divided into retrieval tasks, which typically involve texts of significantly different lengths, and STS/summarization tasks, which involve texts of similar lengths. We examine how Length Collapse affects model performance across these different task types.
Impact on Classification and Clustering Tasks. As shown in Figure 3 (Left), Length Collapse causes long-text embeddings to cluster at the center, while short-text embeddings spread around the periphery. In KNN classification, this brings clustering centers closer to long texts, resulting in a biased influence and reducing performance.
Impact on Retrieval Tasks. As shown in Figure 3 (Middle), long documents are clustered in a smaller central space, while shorter queries are more dispersed, with richer contextual representations Ethayarajh (2019). While centrally-positioned long documents have higher similarity with all embeddings, their representational space is more limited. This can lead to shorter noise documents (Green Square) appearing more relevant to the query (Purple Circle) due to their broader representational space. Analysis of the NFCorpus dataset, comparing rankings of the top 10% longest and shortest relevant documents (Figure 4), shows that short documents follow an inverted U-shaped distribution, whereas long documents exhibit a more uniform distribution. This is because shorter documents have a larger representational space, making it easier for them to appear at the beginning or end of the ranking list.
(a) ANCE
(b) GIST
(c) E5-4K
Figure 4: The count of relevant documents at different ranking positions, where a smaller ranking position indicates a document is more relevant to the query.
Impact on STS Tasks. As shown in Figure 3 (Right), Length Collapse causes long-text embeddings to cluster in a narrower space, where unrelated embeddings may show higher similarity than relevant pairs. Noise sentences (Green Square) in the long-text space may appear more similar than related sentences (Purple Square). In contrast, short-text embeddings are more spread out, resulting in better performance due to greater separation between noise and relevant sentences.
@@ 4 Mitigating Length Collapse via Temperature Scaling
As discussed in § 2.2, the self-attention matrix applies stronger low-pass filtering to longer texts, resulting in distributional differences between long and short text embeddings, which subsequently leads to a decline in long-text performance across various tasks in the MTEB benchmark. To address this issue and validate our analysis, a straightforward solution is to reduce the filtering rate difference between long and short texts, thereby mitigating the disparity in their embeddings. To achieve this, we propose a scaling technique called Temperature Scaling (TempScale), which directly adjusts the attention map by multiplying $\tau_{s}$ by a constant temperature $\tau$ less than 1, thereby reducing the difference in filtering rates $\sigma_{a}$ between long and short texts.
Figure 5: Two extreme cases of TempScale. The darker the color, the higher the attention score.
Specifically, a larger $\sigma_{s}$ will result in a smaller factor $2\sqrt{1+\frac{1}{e^{2\sigma^{2}_{s}}}}$ in Eqn. 2, which means that as the text length $n$ increases, $\sigma_{a}$ will decrease more slowly, thereby reducing the difference in filtering rates between long and short texts. Therefore, by increasing $\sigma_{s}$, or equivalently decreasing $\tau_{s}$ based on Eqn. 3, we can mitigate the distributional difference between long and short text embeddings.
Formally, let $\bm{A}=\mathrm{softmax}\left(\frac{\bm{X}\bm{W}_{Q}(\bm{X}\bm{W}_{K})^{\top}}{\sqrt{d}}\right)$ denote a self-attention matrix. To decrease the $\tau_{s}$ of $\bm{A}$ based on Eqn. 4, we apply a temperature coefficient $\tau$ to the logits before performing the softmax operation. Specifically, for each row $\bm{p}_{i}$ in the attention score matrix $\frac{\bm{X}\bm{W}_{Q}(\bm{X}\bm{W}_{K})^{\top}}{\sqrt{d}}$, we compute the scaled logits by dividing by a temperature $\tau\in(0,1]$, and then apply the softmax function to obtain the attention weights:
@@
where a lower $\tau$ results a smaller $\tau_{s}$.
Another Intuitive Explanation. We illustrate the effects of temperature scaling using two extreme cases to highlight how TempScale works from a different perspective. As shown in Figure 5, when scaling the matrix $\bm{A}$ with a relatively large $\tau$, the elements of $\bm{A}$ become nearly equal, causing the matrix to filter out all high-frequency information and resulting in identical token embeddings. In contrast, when scaling with a smaller temperature, $\bm{A}$ no longer acts as a weighted sum of token representations, but selects a specific token’s representation, preserving more high-frequency information. Viewing the attention matrix as an adjacency matrix, a higher temperature leads to a denser graph, enhancing information exchange between nodes but losing high-frequency details Oono and Suzuki (2020); Cai and Wang (2020). A lower temperature, on the other hand, creates a sparser graph, preserving more high-frequency information and preventing over-smoothing between nodes.
@@ 5 Experiments
In this section, we first conduct experiments to validate the effectiveness of our TempScale on MTEB and LongEmbed. Then we analyze how different tasks can benefit from TempScale to validate our theoretical analysis.
@@
ANCE 55.27 33.04 29.58 66.32 26.95 49.09 34.02 42.04
+TempScale 55.44 33.28 29.59 66.47 27.39 49.25 34.07 42.21
Relative Improv. (%) 0.29 0.73 0.06 0.22 1.63 0.32 0.17 0.49
GTR 55.10 38.65 29.67 70.11 33.00 54.23 37.33 45.44
+TempScale 55.59 39.52 29.83 70.26 33.56 54.22 37.33 45.76
Relative Improv. (%) 0.88 2.26 0.54 0.21 1.69 -0.01 0.01 0.80
GIST 64.75 44.77 31.14 75.61 37.55 58.55 38.21 50.08
+TempScale 65.12 44.66 32.17 75.61 37.82 58.60 38.35 50.33
Relative Improv. (%) 0.56 -0.25 3.31 -0.01 0.71 0.08 0.36 0.68
BGE 64.79 45.80 31.03 75.88 38.14 58.87 37.46 50.28
+TempScale 64.91 45.79 31.87 75.68 38.46 58.97 38.73 50.63
Relative Improv. (%) 0.19 -0.01 2.71 -0.26 0.84 0.17 3.41 1.01
@@
E5 61.72 38.82 30.58 71.77 29.77 53.12 56.01 48.83
+TempScale 62.15 40.62 31.26 72.17 30.28 53.47 56.88 49.55
Relative Improv. (%) 0.7 4.62 2.25 0.55 1.69 0.65 1.56 1.72
Avg Improv. (%) 0.53 1.47 1.77 0.14 1.31 0.24 1.10 0.94
Table 1: Average of the main metric (see Appendix F) per task on MTEB English subsets and LongEmbd. Relative Improv. means percentage increase over the performance without TempScale and improvements are highlighted with while decreasing values are denoted by .
5.1 TempScale Benefits Embedding Models
Experiment Settings. We evaluate embedding models on long-context retrieval using 4 real-world tasks from LongEmbed Zhu et al. (2024), focusing on factuality in both short-query and long-document settings. For other tasks, we select 39 datasets from MTEB Muennighoff et al. (2023), covering classification, clustering, summarization, STS, retrieval, and reranking. Specifically, for classification, we only downloaded 10 datasets due to open-source restrictions. For retrieval, we select the $5$ datasets with the longest documents to better assess Length Collapse. To comprehensively evaluate TempScale, we select several representative PLM-based embedding models, including (1) ANCE Xiong et al. (2021); (2) GTR Ni et al. (2022); (3) GIST Solatorio (2024); (4) BGE Xiao et al. (2023); (5) E5 Zhu et al. (2024). These models are fine-tuned from various pretrain language models, including BERT Kenton and Toutanova (2019), RoBERTa Liu et al. (2019), and T5 Chung et al. (2024). More descriptions of the datasets and models can be found in Appendix G. When evaluating the embedding models, we set the same $\tau$ on the softmax function for the attention modules across all layers within the range of $\{0.1,0.5,0.6,0.7,0.8,0.9,1.0\}$. The metrics used for different tasks are consistent with MTEB and can be found in Appendix F.
Results. We select the optimal temperature $\tau$ for each model based on their performance across a single task and present the results in Table 1. The findings demonstrate that our proposed method, TempScale, improves performance across a range of general tasks, with an average increase of 0.94%. For long-context retrieval tasks, the average improvement is 1.10%. Text length varies widely across tasks—for example, STS uses very short texts, whereas LongEmbed includes much longer ones (1000+ tokens). TempScale proves effective for both preventing long-text collapse and enhancing short-text embeddings, which improves downstream task performance. Additionally, larger context windows yield greater performance gains, with E5 showing the highest improvement of 1.72%. This may be due to larger windows offering more data for adjustment. Beyond the excellent performance of TempScale, more importantly, it validates the correctness of our theoretical analysis.
@@ 5.2 Further Analysis
How do the Classification and Clustering Tasks Benefit from TempScale? In § 3, we attribute the performance decline in classification and clustering tasks to the distributional differences between long and short texts, which lead the model to assign greater weight to long text embeddings during classifier training. TempScale addresses this by adjusting the longer texts to align with the same space as short texts, as described in Figure 1(c), ensuring that both contribute equally during training. More experiments can be found in Appendix C.6.
How do Retrieval Tasks Benefit from TempScale? In § 3, we attribute the performance decline in retrieval tasks to the fact that short texts have more contextual embeddings. Therefore, TempScale improves the performance of long texts by adjusting the temperature to increase high-frequency information, enabling a more contextual distribution. To validate this, we select NFCorpus and SciFact to investigate the impact. As shown in Table 2, we record the average ranking position of the longest positive documents after applying TempScale at different temperatures, where a lower value indicates the documents are ranked higher. The experimental results show that a lower temperature causes the model to rank relevant long documents higher. Our method enables long texts to become more contextual, thereby reducing the bias introduced by length. In summary, TempScale enhances the performance of long documents by introducing TempScale. Similar phenomena on more models and datasets are provided in Appendix C.7.
@@
(b) SciFact
How do the STS Tasks Benefit from TempScale? TempScale enhances performance by giving long sentence embeddings the same spatial representation as short texts. To verify this, we record the cosine similarity between random sentences, related sentences, and unrelated sentences at different temperatures, as shown in Table 3. The results show that as the $\tau$ in TempScale decreases, the similarity between related sentences remains relatively high, while the similarity between unrelated sentences is further reduced. Specifically, although the similarity between random sentences increases with $\tau$ in STS13 and STS14, the increase in the unrelated part is not as large as that of the random sentences. This indicates that the distance between unrelated sentences is increasing.
Can the Temperature be Set Based on the Length of the Texts? As shown in Table 2, the ranking of relevant long documents in retrieval tasks improves as the temperature decreases. This indicates that using a smaller temperature $\tau$ for long texts can better align the embeddings of long and short texts within the same distribution. For longer texts, better performance can be achieved by setting a smaller $\tau$, and the discussions can be found in Appendix C.7.
Temperature STS12 STS13 STS14
Random Related Unrelated Random Related Unrelated Random Related Unrelated
1.0 0.9097 0.9611 0.7973 0.8707 0.9470 0.7931 0.8877 0.9494 0.7889
0.9 0.9096 0.9612 0.7937 0.8721 0.9481 0.7942 0.8886 0.9499 0.7893
0.8 0.9097 0.9612 0.7926 0.8741 0.9490 0.7969 0.8897 0.9501 0.7912
Table 3: Average cosine similarities across different temperature settings.